#1) What percent of 25 is 5?
#2) 25 is what percent of 5?
Please help, the wordings are confusing me

Answer by fcabanski(1391)

The key word is "of". Of means multiply. What % of 25 is 5 means x%*25=5 or 5*x%=25.


x% = 5/25 = .2. Remember % means a number divided by 100, so multiply .2 by 100 to find the %. It's .2 * 100 = 20%.


For the second one it's 25 = x%*5. So x% = 25/5 = 5*100 = 500%.

Hope the solution helped. Sometimes you need more than a solution. Contact fcabanski@hotmail.com for online, private tutoring, or personalized problem solving (quick for groups of problems.)